### v1

# Some Plancherel identities for unbounded subsets of $\mathbb{R}$ in duality

Abstract.
In relation to Fuglede’s conjecture, we establish several Plancherel-type identities and demonstrate the surjectivity of the Fourier transform between certain unbounded tiling sets of $\mathbb{R}$ that are in duality. In the terminology commonly used in the context of Fuglede’s conjecture, our result states that an open set tiles $\mathbb{R}$ by the finite set $\{0,1,\dots,p-1\}$ if and only if it admits a spectrum (or, equivalently, a dual pair measure) given by the Lebesgue measure on $\left[-\tfrac{1}{2p},\tfrac{1}{2p}\right]+\mathbb{Z}$.
Key words and phrases:
differential operator, self-adjoint operator, Fourier bases, Fourier transform, Plancherel identity, Fuglede conjecture, lattice, tile
2010 Mathematics Subject Classification:
47E05,42A16
1. Introduction
In 1958, Irving Segal posed the following question to Bent Fuglede: Let $\Omega$ be an open subset of $\mathbb{R}^{d}$. Consider the partial differential operators
$D_{1}=\frac{1}{2\pi i}\frac{\partial}{\partial x_{1}},\dots,D_{d}=\frac{1}{2\pi i}\frac{\partial}{\partial x_{d}},$
defined on the space $C_{0}^{\infty}(\Omega)$ of smooth, compactly supported functions on $\Omega$. Under what conditions do these differential operators admit commuting (unbounded) self-adjoint extensions $H_{1},\dots,H_{d}$ on $L^{2}(\Omega)$, where commutation is understood in the sense of their spectral measures?
In his seminal 1974 paper [Fug74], Fuglede provided an answer to Segal’s question in the case where $\Omega$ is a connected, finite-measure, Nikodym domain—that is, when the Poincaré inequality holds:
Theorem 1.1.
[Fug74, Theorem I] Let $\Omega\subset\mathbb{R}^{d}$ be an open, connected Nikodym region of finite measure. Define
$e_{\lambda}(x)=e^{2\pi i\lambda\cdot x},\quad(x\in\mathbb{R}^{d},\lambda\in\mathbb{R}^{d}).$
There exist commuting self-adjoint extensions of the differential operators $\{D_{j}:j=1,\dots,d\}$ if and only if there exists a subset $\Lambda$ of $\mathbb{R}^{d}$ such that the family of exponential functions
$\{e_{\lambda}:\lambda\in\Lambda\}$
forms an orthogonal basis for $L^{2}(\Omega)$.
Definition 1.2.
A measurable set $\Omega\subset\mathbb{R}^{d}$ of finite Lebesgue measure is called spectral if there exists a set $\Lambda\subset\mathbb{R}^{d}$ such that the family of exponential functions
$\{e_{\lambda}:\lambda\in\Lambda\}$
forms an orthogonal basis for $L^{2}(\Omega)$. In this case, $\Lambda$ is called a spectrum for $\Omega$.
We say that $\Omega$ tiles $\mathbb{R}^{d}$ by translations if there exists a set $\mathcal{T}\subset\mathbb{R}^{d}$ such that the translates $\{\Omega+t:t\in\mathcal{T}\}$ form a partition of $\mathbb{R}^{d}$ up to measure zero. The set $\mathcal{T}$ is then called a tiling set for $\Omega$, and we say that $\Omega$ tiles $\mathbb{R}^{d}$ by $\mathcal{T}$.
Thus, in this framework, when $\Omega$ is connected, has finite measure, and is a Nikodym domain, there exist commuting self-adjoint extensions of the differential operators $\{D_{j}\}$ if and only if $\Omega$ is spectral. Because this characterization is somewhat abstract, Fuglede proposed his famous conjecture:
Conjecture 1.3.
[Fuglede’s Conjecture] A measurable subset $\Omega\subset\mathbb{R}^{d}$ of finite Lebesgue measure is spectral if and only if it tiles $\mathbb{R}^{d}$ by translations.
The conjecture was later shown to be false in dimensions $d\geq 3$ [Tao04, FMM06], but it remains true for convex domains [LM22].
In [Ped87], Steen Pedersen extended Fuglede’s result by removing both the Nikodym condition and the finite-measure assumption. Naturally, the definition of a spectral set had to be adapted for infinite-measure domains (since $e_{\lambda}$ is not square-integrable in that case).
Definition 1.4.
For a function $f\in L^{1}(\mathbb{R}^{d})$, define its classical Fourier transform as
$\hat{f}(x)=\int_{\mathbb{R}^{d}}f(x)e^{-2\pi it\cdot x}\,dx,\quad(t\in\mathbb{R}^{d}).$
Let $\Omega\subset\mathbb{R}^{d}$ be measurable and let $\mu$ be a positive Radon measure on $\mathbb{R}^{d}$. We say that $(\Omega,\mu)$ is a spectral pair if: (1) for each $f\in L^{1}(\Omega)\cap L^{2}(\Omega)$, the continuous function $t\mapsto\hat{f}(t)$ satisfies $\int|\hat{f}|^{2}\,d\mu<\infty$; and (2) the mapping $f\mapsto\hat{f}$ from $L^{1}(\Omega)\cap L^{2}(\Omega)\subset L^{2}(\Omega)$ into $L^{2}(\mu)$ is isometric and has dense range.
This map then extends by continuity to an isometric isomorphism
$\mathscr{F}:L^{2}(\Omega)\to L^{2}(\mu).$
The isometry property means that we have a Plancherel identity implemented by taking the Fourier transform of functions in $L^{2}(\Omega)$ and restricting it to the support of the measure $\mu$ and integrating the absolute value squared of this restriction of the Fourier transform against the measure $\mu$.
The set $\Omega$ is called spectral if there exists a measure $\mu$ such that $(\Omega,\mu)$ forms a spectral pair; $\mu$ is then called a pair measure or a dual measure for $\Omega$. When $\Omega$ has finite measure, this definition coincides with the earlier one [Ped87, Corollary 1.11], and the dual measure $\mu$ is the counting measure on the spectrum $\Lambda$.
With this broader definition, Pedersen generalized Fuglede’s theorem by eliminating both the Nikodym and finite-measure conditions, while maintaining the assumption that $\Omega$ is connected.
Theorem 1.5.
[Ped87, Theorem 2.2] Let $\Omega$ be an open, connected subset of $\mathbb{R}^{d}$. Then the partial differential operators $\{D_{j}\}$ admit commuting self-adjoint extensions if and only if $\Omega$ is a spectral set.
The results were then extended to disconnected domains in [CD25]. In this case, if $\Omega$ is spectral, then there are commuting self-adjoint extensions of the differential operators $\{D_{j}\}$. However, the converse has to be adjusted a little: the Fourier transform in Definition 1.4 is allowed to have some weights which differ on the components of $\Omega$, and the resulting condition is necessary and sufficient.
In the 1974 paper Fuglede proved that his conjecture is true when the spectrum or the tiling set is a lattice in $\mathbb{R}^{d}$.
Theorem 1.6.
Let $\Omega$ be a measurable subset of $\mathbb{R}^{d}$ of finite measure and let $A$ be some invertible $d\times d$ matrix. Then $\Omega$ tiles $\mathbb{R}^{d}$ by the lattice $A\mathbb{Z}^{d}$ if and only if $\Omega$ has spectrum the dual lattice $(A^{T})^{-1}\mathbb{Z}^{d}$.
Many examples of spectral sets are known in the case when $\Omega$ has finite measure, and the spectrum can be a non-lattice, see e.g, [Kol24]. However, much less is known for sets of infinite measure. Of course, the simplest, and most important example of a spectral set of infinite measure is $\mathbb{R}^{d}$, with pair measure the Lebesgue measure on $\mathbb{R}^{d}$. To obtain some other examples, a result similar to Theorem 1.6 was proved for sublattices of $\mathbb{R}^{d}$ in [CD25]:
Theorem 1.7.
Write $\mathbb{R}^{d}=\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}$. Let $A$ be an invertible real $d\times d$ matrix. The set $\Omega$ tiles with the discrete subgroup $\mathcal{T}=A(\mathbb{Z}^{d_{1}}\times\{0\})$ if and only if $\Omega$ has pair measure $A^{T}({}_{\mathbb{Z}^{d_{1}}}\times\mathfrak{m}_{d_{2}})$ which is supported on the dual set $\mathcal{T}^{*}$,
[…]
Here the Dirac comb $_{\mathbb{Z}^{d_{1}}}$ is the counting measure on $\mathbb{Z}^{d_{1}}$, and $\mathfrak{m}_{d_{2}}$ is the Lebesgue measure on $\mathbb{R}^{d_{2}}$.
For a Borel measure $\mu$ on $\mathbb{R}^{d}$, the measure $A^{T}\mu$ is defined by $A^{T}\mu(E)=|\det(A^{T})^{-1}|\mu(A^{T}E)$ for any Borel subset $E$ in $\mathbb{R}^{d}$. Equivalently, for any compactly supported continuous function $f$ on $\mathbb{R}^{d}$:
(1.1) $\int f\,dA^{T}\mu=|\det(A^{T})^{-1}|\int f((A^{T})^{-1}x)\,d\mu(x).$
Note that if the measure $\mu$ is supported on a set $\Lambda$, then $A^{T}\mu$ is supported on $(A^{T})^{-1}\Lambda$.
In this paper we provide a new class of examples of unbounded spectral sets in $\mathbb{R}$ and we prove the following theorem:
Theorem 1.8.
Let $\Omega$ be an open subset in $\mathbb{R}$ and let $p\in\mathbb{N}$, $p\geq 2$. Then $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$ if and only if $\Omega$ is spectral set with pair measure $\mu$ equal to a renormalized Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, $\mu=p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$ .
2. Proof of Theorem 1.8
We will need some definitions and notations.
Definition 2.1.
Let $\Omega_{0}$ and $\Omega$ be two measurable subsets in $\mathbb{R}$, and let $\mathcal{T}$ be a finite or countable subset of $\mathbb{R}$. We say that $\Omega_{0}$ tiles $\Omega$ by $\mathcal{T}$ if $\{\Omega_{0}+t:t\in\mathcal{T}\}$ is a partition of $\Omega$, up to measure zero.
We denote by $T_{a}$ the translation operator on functions on $\mathbb{R}$
$(T_{a})f(x)=f(x-a),\quad(x\in\mathbb{R},a\in\mathbb{R}).$
Definition 2.2.
Let $E$ and $F$ be two measurable subsets of $\mathbb{R}$ and let $p\in\mathbb{N}$. We say that $E$ and $F$ are congruent $\operatorname{mod}p\mathbb{Z}$ if there exists a partition (up to measure zero) of $E$ into possibly empty sets $\{E_{k}:k\in\mathbb{Z}\}$ such that $\{E_{k}+k:k\in\mathbb{Z}\}$ is a partition of $F$ (up to measure zero).
Definition 2.3.
Let $\mu$ be a finite Radon measure on $\mathbb{R}$ and $\Lambda\subset\mathbb{R}$. We say that $\mu$ is a spectral measure with spectrum $\Lambda$ if the family $\{e_{\lambda}:\lambda\in\Lambda\}$ forms an orthogonal basis for $L^{2}(\mu)$.
Definition 2.4.
For a measurable subset $E$ of $\mathbb{R}$, we denote by $|E|$ its Lebesgue measure and by $\chi_{E}$ its characteristic function. For a finite subset $F$, we denote by $|F|$ its cardinality. We denote by $\delta_{\gamma}$ the Dirac measure at the point $\gamma$ in $\mathbb{R}$. For a discrete subset $\Gamma$ of $\mathbb{R}$, we define the Dirac comb on $\Gamma$ to be the counting measure on $\Gamma$,
$_{\Gamma}=\sum_{\gamma\in\Gamma}\delta_{\gamma}.$
For a measurable subset $E$ of $\mathbb{R}$, we denote by $\mbox{Leb}_{E}$ the Lebesgue measure restricted to $E$.
For a function $f$ on $\mathbb{R}$ we denote its periodization by
$\operatorname*{Per}(f)(t)=\sum_{k\in\mathbb{Z}}f(t+k),\quad(t\in\mathbb{R}).$
We begin now the proof of our main result.
Proof of Theorem 1.8.
We assume that $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$.
Proposition 2.5.
Suppose $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$. Define $\Omega_{0}=\Omega\cap[0,p)$.
  1. (i)
    $\Omega+kp\subset\Omega$ for all $k\in\mathbb{Z}$.
  2. (ii)
    $\Omega_{0}$ tiles $\Omega$ by $p\mathbb{Z}$; in particular
    $\Omega=\Omega_{0}+p\mathbb{Z}.$
  3. (iii)
    $\Omega_{0}$ tiles $\mathbb{R}$ by $\mathbb{Z}$. Therefore $\Omega_{0}$ has measure 1, has spectrum $\mathbb{Z}$, and is congruent to $[0,1)$ modulo $\mathbb{Z}$.
Proof.
Let $x\in\Omega$. Then $x+p\in\Omega$. Indeed, since $\Omega$ tiles with $\{0,\dots,p-1\}$ we can write $x+p=y+j$ for $y\in\Omega$ and $j\in\{0,\dots,p-1\}$. If $j\neq 0$, then x+(pj)=yx+(p-j)=y so $y\in\Omega\cap(\Omega+(p-j))$, which can happen only for $y$ in a measure zero set.
If xp=y+jx-p=y+j with $x\in\Omega$ and $j\in\{0,1,\dots,p-1\}$ then $\Omega\ni x=y+p+j\in\Omega+j$ and therefore $j$ has to be 0 and $x-p\in\Omega$. By induction, we get $x+pk\in\Omega$, for $k\in\mathbb{Z}$. This proves (i).
For (ii), it is clear from (i) that $\Omega_{0}+p\mathbb{Z}\subseteq\Omega$. Now take $x\in\Omega$. Then we can write $x=y+pk$ for some $y\in[0,p)$ and $k\in\mathbb{Z}$. With (i), $y=x-kp\in\Omega$. Thus $y\in\Omega_{0}$ and $x\in\Omega_{0}+p\mathbb{Z}$. This proves the reverse inclusion.
Since $\Omega_{0}$ is contained in $[0,p)$, the sets $\Omega_{0}+pk$, $k\in\mathbb{Z}$, are mutually disjoint. This proves (ii).
For (iii), note that
$\Omega_{0}+\mathbb{Z}=\Omega_{0}+p\mathbb{Z}+\{0,\dots,p-1\}=\Omega+\{0,\dots,p-1\}=\mathbb{R}.$
The sets $\Omega_{0}+pk+j$, with $k\in\mathbb{Z}$ and $j\in\{0,\dots,p-1\}$ are disjoint by (ii) and the fact that $\Omega$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$. Thus $\Omega_{0}$ tiles $\mathbb{R}$ by $\mathbb{Z}$.
With Theorem 1.6, we get that $\Omega$ has spectrum $\mathbb{Z}$. Also, let $\Omega_{0,k}=\Omega_{0}\cap[k,k+1)$ and let $I_{k}=\Omega_{0,k}-k\subset[0,1)$, for all $k\in\mathbb{Z}$. Then the sets $\{I_{k}:k\in\mathbb{Z}\}$ are disjoint (otherwise the sets $\Omega_{0}+k$ are not), and their union is $[0,1)$ because, for $x\in[0,1)$ there is a unique $k\in\mathbb{Z}$ and $x_{0}\in\Omega_{0}$ such that $x=x_{0}-k$. So $x_{0}\in\Omega_{0,k}$ and $x\in I_{k}$. Thus $\Omega_{0}$ is congruent to $[0,1)$ modulo $\mathbb{Z}$ and
$|\Omega_{0}|=\sum_{k\in\mathbb{Z}}|\Omega_{0,k}|=\sum_{k\in\mathbb{Z}}|I_{k}|=1.$
Proposition 2.6.
For $n\in\mathbb{N}$ let
$\Omega_{n}:=\Omega_{0}+p\{-n,-n+1,\dots,n-1,n\}.$
  1. (i)
    The sets $\Omega_{n}$ are increasing and cover $\Omega$:
    $\cup_{n\in\mathbb{N}}\Omega_{n}=\Omega.$
  2. (ii)
    The set $\Omega_{n}$ tiles $\Omega$ by $p(2n+1)\mathbb{Z}$.
  3. (iii)
    The set $\Omega_{n}$ has spectrum
    $\Lambda_{n}:=\frac{1}{p(2n+1)}\left\{-n,-n+1,\dots,n\right\}+\mathbb{Z}.$
Proof.
(i) is clear, from Proposition 2.5(ii). For (ii), note that the set $L_{n}=\{-n,\dots,n\}$ tiles $\mathbb{Z}$ by $(2n+1)\mathbb{Z}$. Therefore
$\Omega_{n}+p(2n+1)\mathbb{Z}=\Omega_{0}+pL_{n}+p(2n+1)\mathbb{Z}=\Omega_{0}+p\mathbb{Z},$
and the sets $\Omega_{n}+p(2n+1)k=\Omega_{0}+pL_{n}+p(2n+1)k$, $k\in\mathbb{Z}$ are disjoint.
For (iii), we will use the following well known lemma (see e.g., [JP98]). We include the proof for convenience.
Lemma 2.7.
Let $\mu$ be finite Radon measure on $\mathbb{R}$ with total measure $\mu(\mathbb{R})=m$, and let $\Lambda$ be a finite or countable subset of $\mathbb{R}$. The measure $\mu$ has spectrum $\Lambda$ if and only if
(2.1) $\sum_{\lambda\in\Lambda}|\hat{\mu}(t+\lambda)|^{2}=m^{2}.$
Proof.
Suppose $\Omega$ has spectrum $\Lambda$. Then
$\{\frac{1}{\sqrt{m}}e_{\lambda}:\lambda\in\Lambda\}$
is an orthonormal basis for $L^{2}(\mu)$. Applying the Parseval identity to the function $e_{-t}$, $t\in\mathbb{R}$ we obtain
$m=\|e_{-t}\|^{2}=\sum_{\lambda\in\Lambda}\left|\left\langle e_{-t}\,,\,\frac{1}{\sqrt{m}}e_{\lambda}\right\rangle_{L^{2}(\mu)}\right|^{2}=\frac{1}{m}\sum_{\lambda\in\Lambda}\left|\hat{\mu}(t+\lambda)\right|^{2}$
and this implies (2.1).
For the converse, if we plug $t=-\lambda_{0}$ in (2.1), since $\hat{\mu}(0)=|\Omega|=m$, we get that $\hat{\mu}(-\lambda_{0}+\lambda)=0$ for $\lambda\neq\lambda_{0}$, and this proves that the exponentials $e_{\lambda}$ and $e_{\lambda_{0}}$ are orthogonal. With the same computation as above, (2.1) then shows that
$\|e_{-t}\|^{2}=\sum_{\lambda\in\Lambda}\left|\left\langle e_{-t}\,,\,\frac{1}{\sqrt{m}}e_{\lambda}\right\rangle_{L^{2}(\mu)}\right|^{2}.$
On the right hand side we have the norm squared of the projection of $e_{-t}$ onto the closed span of the functions $\{e_{\lambda}:\lambda\in\Lambda\}$. This shows that $e_{-t}$ is in this closed span. By the Stone-Weierstrass theorem and by approximation by continuous compactly supported functions, the functions $e_{-t}$ span the entire space $L^{2}(\mu)$. It follows that $\{e_{\lambda}:\lambda\in\Lambda\}$ also span the entire space $L^{2}(\mu)$ and therefore they form an orthogonal basis for it. ∎
Returning to the proof of Proposition 2.6(iii), we apply Lemma 2.7 to the Lebesgue measure on $\Omega_{n}$. Note that this measure is the convolution $\mbox{Leb}_{\Omega_{n}}=\mbox{Leb}_{\Omega_{0}}\ast{}_{pL_{n}}$. Therefore its Fourier transform is
$\hat{\chi}_{\Omega_{n}}(t)=\hat{\chi}_{\Omega_{0}}(t)\hat{\Sha}_{pL_{n}}(t).$
We also have
$\hat{\Sha}_{pL_{n}}(t)=\int e^{-2\pi itx}\,d{}_{pL_{n}}(x)=\sum_{l=-n}^{n}e^{-2\pi iplt},$
which implies that $\hat{\Sha}_{pL_{n}}$ has period $1/p$.
It is also easy to check that the measure $_{pL_{n}}$ has spectrum $\frac{1}{p(2n+1)}L_{n}$ (this is basically the discrete Fourier transform on $\mathbb{Z}_{2n+1}$, rescaled by $p$). Therefore, with Lemma 2.7,
(2.2) $\sum_{j\in L_{n}}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}\right)\right|^{2}=|L_{n}|^{2}=(2n+1)^{2},\quad(t\in\mathbb{R}).$
Since $\Omega_{0}$ has spectrum $\mathbb{Z}$, we have that
(2.3) $\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{0}}(t+k)\right|^{2}=|\Omega_{0}|^{2}=1,\quad(t\in\mathbb{R}).$
We compute
$\sum_{j\in L_{n}}\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{n}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}$
$\sum_{j\in L_{n}}\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{0}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}$
and using the periodicity of $\hat{\Sha}_{pL_{n}}$,
$\sum_{j\in L_{n}}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}\right)\right|^{2}\sum_{k\in\mathbb{Z}}\left|\hat{\chi}_{\Omega_{0}}\left(t+\frac{j}{p(2n+1)}+k\right)\right|^{2}$
and with (2.3), and then (2.2),
$\sum_{j\in L_{n}}\left|\hat{\Sha}_{pL_{n}}\left(t+\frac{j}{p(2n+1)}\right)\right|^{2}\cdot 1=(2n+1)^{2}=|\Omega_{n}|^{2}.$
Then, with Lemma 2.7, we conclude that $\frac{1}{p(2n+1)}L_{n}+\mathbb{Z}$ is a spectrum for $\Omega_{n}$. ∎
Next we prove the first condition for $\Omega$ to be spectral, from Definition 1.4, namely that the Fourier transform is isometric from $L^{2}(\Omega)$ to the $L^{2}$-space of the pair measure, which is $p$ times the Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$. The notation for this $L^{2}$-space is $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx\right)$.
Proposition 2.8.
The Fourier transform
$L^{2}(\Omega)\ni f\to\hat{f}|_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}\in L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx\right)$
is an isometry.
Proof.
Let $f$ be a $C^{\infty}$-function compactly supported in $\Omega$. Then $\hat{f}$ is a Schwartz function.
Let $n$ be big enough so that the support of $f$ is covered by $\Omega_{n}=\Omega_{0}+p\{-n,\dots,n\}$. Since $\Omega_{n}$ has spectrum $\frac{1}{p(2n+1)}L_{n}+\mathbb{Z}$, we have, using the Parseval identity,
$\int_{\Omega}|f(x)|^{2}\,dx=\int_{\Omega_{n}}|f(x)|^{2}\,dx=\sum_{j\in L_{n},k\in\mathbb{Z}}\left|\left\langle f\,,\,\frac{1}{\sqrt{2n+1}}e_{\frac{j}{p(2n+1)}+k}\right\rangle_{L^{2}(\Omega_{n})}\right|^{2}$
$\frac{1}{2n+1}\sum_{j\in L_{n},k\in\mathbb{Z}}\left|\hat{f}\left(\frac{j}{p(2n+1)}+k\right)\right|^{2}.$
This is a Riemann sum for the function $|\hat{f}|^{2}$ with sampling points $\frac{j}{p(2n+1)}+k$, $j\in L_{n}$, $k\in\mathbb{Z}$, and thus the length of the intervals of the partition is $\frac{1}{p(2n+1)}$. Since the common factor is $\frac{1}{2n+1}$ we have to adjust it by the constant $p$, and therefore the function that is integrated is actually $p|\hat{f}|^{2}$ on the set $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$.
Since $\hat{f}$ is a Schwartz function, these Riemann sums converge to the integral
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}|\hat{f}(t)|^{2}p\,dt.$
This implies that
$\int_{\Omega}|f(x)|^{2}\,dx=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}|\hat{f}(t)|^{2}p\,dt.$
Since $C_{0}^{\infty}(\Omega)$ is dense in $L^{2}(\Omega)$, we get that the Fourier transform establishes an isometry between the two $L^{2}$-spaces.
Proposition 2.9.
The Fourier transform from $L^{2}(\Omega)$ to $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx\right)$ is onto.
Proof.
First, note that since $\Omega_{0}$ is congruent to $[0,1]$ $\operatorname{mod}\mathbb{Z}$, we have for $k\in\mathbb{Z}$,
(2.4) $\hat{\chi}_{\Omega_{0}}(k)=\int_{\Omega_{0}}e^{-2\pi ikx}\,dx=\int_{[0,1]}e^{-2\pi ikx}\,dx=\left\{\begin{array}[]{cc}1,&k=0,\\ 0,&k\neq 0.\end{array}\right$.
The function $\chi_{\Omega}=\chi_{\Omega_{0}+p\mathbb{Z}}$ has period $p$. Therefore, its Fourier transform, as a tempered distribution, is
(2.5) $\hat{\chi}_{\Omega}=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}.$
This follows from the next lemma:
Lemma 2.10.
Let $\tilde{h}$ be a function of period $p$ on $\mathbb{R}$ and $h=\tilde{h}|_{[0,p]}\in L^{2}[0,p]$. Represent $\tilde{h}$ in the Fourier basis $\frac{1}{\sqrt{p}}e_{k/p}$, $k\in\mathbb{Z}$:
$\tilde{h}=\sum_{k\in\mathbb{Z}}\frac{1}{p}\left\langle h\,,\,e_{k/p}\right\rangle_{L^{2}[0,p]}e_{k/p}=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)e_{k/p}.$
Then the Fourier transform of $\tilde{h}$ as a tempered distribution is
(2.6) $\hat{\tilde{h}}=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)\delta_{k/p}.$
Proof.
Let $\Lambda_{\tilde{h}}$ be the tempered distribution associated to $\tilde{h}$,
$\Lambda_{\tilde{h}}(\varphi)=\int_{\mathbb{R}}\tilde{h}(x)\varphi(x)\,dx,\quad(\varphi\mbox{ Schwartz function}).$
Then, for a Schwartz function $\varphi$,
$\hat{\Lambda}_{\tilde{h}}(\varphi)=\Lambda_{\tilde{h}}(\hat{\varphi})=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)\int_{\mathbb{R}}e^{2\pi itk/p}\hat{\varphi}(t)\,dt$
with the Fourier Inversion Formula,
$\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{h}(k/p)\varphi(k/p),$
and (2.6) follows. The sums are convergent because $\sum|\hat{h}(k/p)|^{2}<\infty$ and $\varphi,\hat{\varphi}$ are Schwartz functions, thus they have rapid decay.
Take now some $C_{0}^{\infty}$-function $g$ supported on one of the intervals $\left[-\frac{1}{2p},\frac{1}{2p}\right]+k_{0}$ for some $k_{0}\in\mathbb{Z}$, and let $\varphi$ be its inverse Fourier transform; it is a Schwartz function. Let
$f=\chi_{\Omega}\cdot\varphi.$
Then
$\hat{f}(t)=\hat{\chi}_{\Omega}\ast\hat{\varphi}(t)=\left(\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}\right)\ast g(t)=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)g(t-k/p).$
We need to restrict $\hat{f}$ to $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$. Note that g(tk/p)g(t-k/p) is supported on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\frac{k}{p}+k_{0}$, and these sets are disjoint for different $k$’s. Thus, when we restrict $\hat{f}$ to $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, we will keep only the terms with $k=pl$ for $l\in\mathbb{Z}$:
$\hat{f}\cdot\chi_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}(t)=\sum_{l\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(pl/p)g(t-pl/p).$
But, by (2.4), $\hat{\chi}_{\Omega_{0}}(l)=0$ unless $l=0$, and therefore the only term left is $g(t)$. Thus
$\hat{f}|_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}=g.$
Since the linear combinations of functions like $g$ are dense in $L^{2}(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},pdx)$ and since we proved that the map is an isometry, it follows that the map is onto. ∎
Propositions 2.8 and 2.9 show the direct implication: if $\Omega$ tiles with $\{0,1,\dots,p-1\}$ then it is spectral with the desired pair measure.
We move on now to the converse and assume that $\Omega$ is spectral with the given pair measure.
Proposition 2.11.
[…]
Proof.
Let $a\in\mathbb{R}$. We have
$\left\langle f\,,\,g\right\rangle_{L^{2}(\Omega)}=\left\langle e_{a}f\,,\,e_{a}g\right\rangle_{L^{2}(\Omega)}=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}\hat{f}(t-a)\overline{\hat{g}}(t-a)p\,dt$
$\sum_{k\in\mathbb{Z}}\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\hat{f}(t-a+k)\overline{\hat{g}}(t-a+k)p\,dt$
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t-a)p\,dt=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t)p\,dt.$
Let $h(t)=\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t)$. Then the function
$a\to\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a}h(t)\,dt$
is constant.
Let $a_{1}=-\frac{1}{2p}-b-\epsilon$, $a_{2}=-\frac{1}{2p}-b$. We have
$0=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a_{1}}h(t)\,dt-\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a_{2}}h(t)\,dt=\int_{[b+\epsilon,b+\frac{1}{p}+\epsilon]}h(t)\,dt-\int_{[b,b+\frac{1}{p}]}h(t)\,dt$
$\int_{[b+\frac{1}{p},b+\frac{1}{p}+\epsilon]}h(t)\,dt-\int_{[b,b+\epsilon]}h(t)\,dt.$
Divide by $\epsilon$ and, with the Lebesgue Differentiation Theorem, we get that for almost every $b\in\mathbb{R}$,
$h(b+1/p)=h(b),$
[…]
Proposition 2.12.
Suppose $\Omega$ is spectral with pair measure $p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$. Then $\Omega\cap(\Omega+j)=\emptyset$ up to measure zero, for all $j\in\mathbb{Z}$, $j\not\equiv 0\operatorname{mod}p$.
Proof.
Let $f,g\in L^{2}(\Omega)$. We compute, with Plancherel’s identity,
$\left\langle f\,,\,T_{j}g\right\rangle_{L^{2}(\mathbb{R})}=\int_{\mathbb{R}}{\hat{f}(t)}\overline{e^{-2\pi ijt}\hat{g}(t)}\,dt=\sum_{l\in\mathbb{Z}}\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\hat{f}(t+l/p)e^{2\pi ij(t+l/p)}\overline{\hat{g}}(t+l/p)\,dt$
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\sum_{i=0}^{p-1}\sum_{m\in\mathbb{Z}}\hat{f}(t+i/p+m)\overline{\hat{g}}(t+i/p+m)e^{2\pi ij(t+i/p)}\,dt$
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\sum_{i=0}^{p-1}e^{2\pi ij(t+i/p)}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t+i/p)\,dt$
with Propositon 2.11,
$\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]}\operatorname*{Per}(\hat{f}\overline{\hat{g}})(t)e^{2\pi ijt}\sum_{i=0}^{p-1}e^{2\pi iji/p}\,dt=0,$
since $j\not\equiv 0(\operatorname{mod}p)$.
Suppose now that $\Omega\cap(\Omega+j)$ has positive measure. Then take $E\subset\Omega\cap(\Omega+j)$ of finite measure. Let $f=\chi_{E}$ and $g=\chi_{E-j}$. Then $f,g\in L^{2}(\Omega)$ and $T_{j}g=f$ Thus $0=\left\langle f\,,\,T_{j}g\right\rangle_{L^{2}(\mathbb{R})}=\|f\|_{L^{2}(\mathbb{R})}^{2}>0$, a contradiction . ∎
We want to show that $\Omega$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$. We will show next that $\Omega$ is contained in a larger set $\tilde{\Omega}$ which tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$.
The set $\Omega^{\prime}=\Omega+p\mathbb{Z}$ is clearly invariant under translations by $p\mathbb{Z}$ and we show that
(2.7) $\{\Omega^{\prime}+j:j\in\{0,1,\dots,p-1\}\}\mbox{ are disjoint sets}.$
If not, there exist $k_{1},k_{2}\in\mathbb{Z}$, $j_{1},j_{2}\in\{0,\dots,p-1\}$ such that $(\Omega+pk_{1}+j_{1})\cap(\Omega+pk_{2}+j_{2})\neq\emptyset$. We can assume $j_{1}>j_{2}$ and then $j_{1}-j_{2}\in\{0,\dots,p-1\}$ and $(\Omega+p(k_{1}-k_{2})+(j_{1}-j_{2}))\cap\Omega\neq\emptyset$, which contradicts Proposition 2.12.
Lemma 2.13.
Define
$\Omega_{0}:=\Omega^{\prime}\cap[0,p).$
  1. (i)
    The sets $\{\Omega_{0}+k:k\in\mathbb{Z}\}$ are disjoint.
  2. (ii)
    $\Omega^{\prime}=\Omega+p\mathbb{Z}=\Omega_{0}+p\mathbb{Z}$.
Proof.
Consider two sets $\Omega_{0}+pl+j$ and $\Omega_{0}+pl^{\prime}+j^{\prime}$ with $l,l^{\prime}\in\mathbb{Z}$ and $j,j^{\prime}\in\{0,\dots,p-1\}$. If $j\neq j^{\prime}$ then the intersection of these two sets in contained in $(\Omega^{\prime}+j)\cap(\Omega^{\prime}+j^{\prime})=\emptyset$. If $j=j^{\prime}$ and the intersection of the two sets is nonempty, then $(\Omega_{0}+pl)\cap(\Omega_{0}+pl^{\prime})\neq\emptyset$ and this is impossible because $\Omega_{0}$ is contained in $[0,p)$. This proves (i).
For (ii), we clearly have $\Omega_{0}+p\mathbb{Z}\subset\Omega+p\mathbb{Z}$. For the reverse inclusion, let $x\in\Omega^{\prime}$. Then, there exists $k\in\mathbb{Z}$ such that $x-pk\in[0,p)$. Since $\Omega^{\prime}$ is invariant under translations by $p\mathbb{Z}$, it follows that $x-pk\in\Omega^{\prime}\cap[0,p)=\Omega_{0}$ and thus $x\in\Omega_{0}+pk$. ∎
Lemma 2.14.
Define $R:=[0,1)\setminus((\Omega_{0}+\mathbb{Z})\cap[0,1))$. $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $\mathbb{Z}$ and $\tilde{\Omega}:=(\Omega_{0}\cup R)+p\mathbb{Z}$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$ and contains $\Omega$.
Proof.
Note that $\Omega_{0}+\mathbb{Z}$ and $R+\mathbb{Z}$ are disjoint. Otherwise, there exist $x_{0}\in\Omega_{0}$, $k\in\mathbb{Z}$, $r\in R$ and $l\in\mathbb{Z}$ such that $x_{0}+k=r+l$. So $r=x_{0}+(k-l)\in(\Omega_{0}+\mathbb{Z})\cap[0,1)$, which contradicts the definition of $R$. The sets $\{R+k:k\in\mathbb{Z}\}$ are disjoint because $R$ is contained in $[0,1)$. The sets $\{\Omega_{0}+k:k\in\mathbb{Z}\}$ are disjoint, by Lemma 2.13. It follows that the sets $\{(\Omega_{0}\cup R)+k:k\in\mathbb{Z}\}$ are disjoint. Also,
$(\Omega_{0}\cup R)+\mathbb{Z}=(\Omega_{0}+\mathbb{Z}+\mathbb{Z})\cup(R+\mathbb{Z})=((\Omega_{0}+\mathbb{Z})\cup R)+\mathbb{Z}\supseteq[0,1)+\mathbb{Z}=\mathbb{R}.$
This shows that $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $[0,1)$.
For $\tilde{\Omega}=(\Omega_{0}\cup R)+p\mathbb{Z}$, since $\Omega_{0}$ tiles $\mathbb{R}$ with $\mathbb{Z}$, we get that the sets $\{\tilde{\Omega}+j=(\Omega_{0}\cup R)+j+p\mathbb{Z}:j\in\{0,\dots,p-1\}\}$ are disjoint, and
$\tilde{\Omega}+\{0,\dots,p-1\}=(\Omega_{0}\cup R)+p\mathbb{Z}+\{0,\dots,p-1\}=(\Omega_{0}\cup R)+\mathbb{Z}=\mathbb{R}.$
Thus, $\tilde{\Omega}$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$.
By Lemma 2.13, $\Omega\subseteq\Omega_{0}+p\mathbb{Z}\subseteq\tilde{\Omega}$.
Lemma 2.15.
$\Omega=\tilde{\Omega}$ and therefore $\Omega$ tiles $\mathbb{R}$ by $\{0,\dots,p-1\}$.
Proof.
Suppose the inclusion $\Omega\subset\tilde{\Omega}$ is proper, so the complement $\tilde{\Omega}\setminus\Omega$ has positive Lebesgue measure.
[…]
$L^{2}(\tilde{\Omega})\ni f\to\hat{f}|_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}=:\mathcal{F}_{r}f\in L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},p\,dx\right)$
is a surjective isometry. Therefore $\mathcal{F}_{r}f\perp\mathcal{F}_{r}g$, and also $\mathcal{F}_{r}g\neq 0$.
On the other hand, by hypothesis, $\mathcal{F}_{r}$ is a surjective isometry from $L^{2}(\Omega)$ to the same range $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},p\,dx\right)$. Thus $\{\mathcal{F}_{r}f:f\in L^{2}(\Omega)\}$ is the entire range $L^{2}\left(\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z},p\,dx\right)$; therefore, $\mathcal{F}_{r}g$ is orthogonal to the entire range, and thus it is orthogonal to itself, so it has to be zero, a contradiction.
In conclusion, $\Omega=\tilde{\Omega}$ and thus $\Omega$ tiles $\mathbb{R}$ with $\{0,\dots,p-1\}$, and this shows the converse implication in our Theorem 1.8. ∎

### v2

@@
 Let $\Omega$ be an open subset in $\mathbb{R}$ and let $p\in\mathbb{N}$, $p\geq 2$.
-Then $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$ if and only if $\Omega$ is spectral set with pair measure $\mu$ equal to a renormalized Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, $\mu=p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$ .
+Then $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$ if and only if $\Omega$ is a spectral set with pair measure $\mu$ equal to a renormalized Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, $\mu=p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$ .
 2. Proof of Theorem 1.8
@@
 Definition 2.2.
-Let $E$ and $F$ be two measurable subsets of $\mathbb{R}$ and let $p\in\mathbb{N}$.
-We say that $E$ and $F$ are congruent $\operatorname{mod}p\mathbb{Z}$ if there exists a partition (up to measure zero) of $E$ into possibly empty sets $\{E_{k}:k\in\mathbb{Z}\}$ such that $\{E_{k}+k:k\in\mathbb{Z}\}$ is a partition of $F$ (up to measure zero).
+Let $E$ and $F$ be two measurable subsets of $\mathbb{R}$.
+We say that $E$ and $F$ are congruent $\operatorname{mod}\mathbb{Z}$ if there exists a partition (up to measure zero) of $E$ into possibly empty sets $\{E_{k}:k\in\mathbb{Z}\}$ such that $\{E_{k}+k:k\in\mathbb{Z}\}$ is a partition of $F$ (up to measure zero).
 Definition 2.3.
@@
 Thus $\Omega_{0}$ tiles $\mathbb{R}$ by $\mathbb{Z}$.
-With Theorem 1.6, we get that $\Omega$ has spectrum $\mathbb{Z}$.
+With Theorem 1.6, we get that $\Omega_{0}$ has spectrum $\mathbb{Z}$.
 Also, let $\Omega_{0,k}=\Omega_{0}\cap[k,k+1)$ and let $I_{k}=\Omega_{0,k}-k\subset[0,1)$, for all $k\in\mathbb{Z}$.
@@
 The measure $\mu$ has spectrum $\Lambda$ if and only if
-(2.1) $\sum_{\lambda\in\Lambda}|\hat{\mu}(t+\lambda)|^{2}=m^{2}.$
+(2.1) $\sum_{\lambda\in\Lambda}|\hat{\mu}(t+\lambda)|^{2}=m^{2},$
+where $\hat{\mu}$ is the Fourier transform of the measure $\mu$,
+$\hat{\mu}(t)=\int e^{-2\pi itx}\,d\mu(x),\quad(t\in\mathbb{R}).$
 Proof.
@@
 The function $\chi_{\Omega}=\chi_{\Omega_{0}+p\mathbb{Z}}$ has period $p$. Therefore, its Fourier transform, as a tempered distribution, is
-(2.5) $\hat{\chi}_{\Omega}=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}.$
+(2.5) $\hat{\chi}_{\Omega}=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}.$
 This follows from the next lemma:
@@
 Let
-$f=\chi_{\Omega}\cdot\varphi.$
+$f=p\chi_{\Omega}\cdot\varphi.$
 Then
-$\hat{f}(t)=\hat{\chi}_{\Omega}\ast\hat{\varphi}(t)=\left(\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}\right)\ast g(t)=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)g(t-k/p).$
+$\hat{f}(t)=p\hat{\chi}_{\Omega}\ast\hat{\varphi}(t)=\left(\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}\right)\ast g(t)=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)g(t-k/p).$
 We need to restrict $\hat{f}$ to $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$.
@@
 is constant.
-Let $a_{1}=-\frac{1}{2p}-b-\epsilon$, $a_{2}=-\frac{1}{2p}-b$. We have
+Let $a_{1}=\frac{1}{2p}+b+\epsilon$, $a_{2}=\frac{1}{2p}+b$, for $b\in\mathbb{R}$ and $\epsilon>0$. We have
 $0=\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a_{1}}h(t)\,dt-\int_{\left[-\frac{1}{2p},\frac{1}{2p}\right]-a_{2}}h(t)\,dt=\int_{[b+\epsilon,b+\frac{1}{p}+\epsilon]}h(t)\,dt-\int_{[b,b+\frac{1}{p}]}h(t)\,dt$
@@
 Consider two sets $\Omega_{0}+pl+j$ and $\Omega_{0}+pl^{\prime}+j^{\prime}$ with $l,l^{\prime}\in\mathbb{Z}$ and $j,j^{\prime}\in\{0,\dots,p-1\}$.
-If $j\neq j^{\prime}$ then the intersection of these two sets in contained in $(\Omega^{\prime}+j)\cap(\Omega^{\prime}+j^{\prime})=\emptyset$.
+If $j\neq j^{\prime}$ then the intersection of these two sets is contained in $(\Omega^{\prime}+j)\cap(\Omega^{\prime}+j^{\prime})=\emptyset$.
 If $j=j^{\prime}$ and the intersection of the two sets is nonempty, then $(\Omega_{0}+pl)\cap(\Omega_{0}+pl^{\prime})\neq\emptyset$ and this is impossible because $\Omega_{0}$ is contained in $[0,p)$.
@@
 Also,
-$(\Omega_{0}\cup R)+\mathbb{Z}=(\Omega_{0}+\mathbb{Z}+\mathbb{Z})\cup(R+\mathbb{Z})=((\Omega_{0}+\mathbb{Z})\cup R)+\mathbb{Z}\supseteq[0,1)+\mathbb{Z}=\mathbb{R}.$
-This shows that $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $[0,1)$.
-For $\tilde{\Omega}=(\Omega_{0}\cup R)+p\mathbb{Z}$, since $\Omega_{0}$ tiles $\mathbb{R}$ with $\mathbb{Z}$, we get that the sets $\{\tilde{\Omega}+j=(\Omega_{0}\cup R)+j+p\mathbb{Z}:j\in\{0,\dots,p-1\}\}$ are disjoint, and
+$(\Omega_{0}\cup R)+\mathbb{Z}=(\Omega_{0}+\mathbb{Z})\cup(R+\mathbb{Z})=((\Omega_{0}+\mathbb{Z})\cup R)+\mathbb{Z}\supseteq[0,1)+\mathbb{Z}=\mathbb{R}.$
+This shows that $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $\mathbb{Z}$.
+For $\tilde{\Omega}=(\Omega_{0}\cup R)+p\mathbb{Z}$, since $\Omega_{0}\cup R$ tiles $\mathbb{R}$ with $\mathbb{Z}$, we get that the sets $\{\tilde{\Omega}+j=(\Omega_{0}\cup R)+j+p\mathbb{Z}:j\in\{0,\dots,p-1\}\}$ are disjoint, and
 $\tilde{\Omega}+\{0,\dots,p-1\}=(\Omega_{0}\cup R)+p\mathbb{Z}+\{0,\dots,p-1\}=(\Omega_{0}\cup R)+\mathbb{Z}=\mathbb{R}.$
@@
 In conclusion, $\Omega=\tilde{\Omega}$ and thus $\Omega$ tiles $\mathbb{R}$ with $\{0,\dots,p-1\}$, and this shows the converse implication in our Theorem 1.8. ∎
+Conflict of interest. On behalf of all authors, the corresponding author states that there is no conflict of interest.
+Data Availability Statement. No new data were created or analyzed in this study. Data sharing is not applicable to this article.
+Funding. No funds, grants, or other support were received during the preparation of this manuscript.
+Author Contribution. All authors contributed equally to the study conception, design, data collection, analysis, and manuscript preparation. Both authors read and approved the final manuscript.